# SLANG: New Concept Comprehension of Large Language Models

### 23 Jan 2024 v1

@@ 3.1.2 Analysis with Linguistic Factors
3.2 Context-based Causal Intervention
Subsequently, we refine our causal intervention approach in the SCM, focusing on the role of context and linguistic factors. To eliminate the entity bias in $W$, it is replaced with a specific placeholder, denoted as $\hat{W}$ (as illustrated in Figure 4 (d)). This alteration aids in isolating the effect of $W$ while excluding entity-specific biases. Further, to concentrate on the contextual content and reduce reliance on shortcuts, $E$ is replaced with $\tilde{E}$, consequently transforming $X$ into $\tilde{X}$. This shift accentuates the role of $R$ in shaping the context, enhancing the model’s capacity to highlight the influence of linguistic factors in a bias-free manner. The updated causal effect formula in the SCM is thus:
$P(Y=y|do(X=x))$
$P(\tilde{W}=\tilde{w}|do(X=x),\tilde{E}=\tilde{e},R=r)$
$P(\tilde{E}=\tilde{e})P(R=r)$
The core idea of FOCUS is to explore how language models can adhere to guidelines to better understand the content of the context. The goal is to enable LLMs to analyze phrases according to usage examples and provide counterfactual interpretations, thereby understanding the evolving semantics of language. To achieve this objective, we propose a four-stage method, here is a description of each stage and its principles:
Direct Inquiry (DI)
The DI stage is the foundational step in our FOCUS methodology. It primarily involves inputting the usage example (context) $X$ and the phrase $W$ directly into the language model. This stage is represented as $Y_{DI}=\mathcal{M}(X,W,E)$, where $\mathcal{M}$ is the function that defines the processing of the language model. The aim here is to assess the direct effect of the context $X$ on the inferred meaning $Y$ in a straightforward and intervention-free manner. This allows the model to naturally derive the meaning within the given context. Special emphasis is placed on maintaining the integrity of all confounders, ensuring that the phrase’s morphology, literal meaning, and context-based interpretation are coherently analyzed. The DI stage sets the baseline for understanding the meaning of the phrase before any further analytical layers are added in the subsequent stages.
Masked Entity Analysis (MEA)
The MEA stage is a crucial component of our FOCUS methodology, designed to mitigate the impact of the phrase’s literal meaning on its interpretation. In this stage, the phrase $W$ is masked within the context $X$, leading to an evaluation of the meaning $Y$ in the absence of direct influence from $W$. This process is denoted as $Y_{MEA},\hat{W}_{MEA}=\mathcal{M}(X_{\text{masked}},E)$, where $\mathcal{M}$ represents the language model. The primary objective is to understand the meaning based on context, prompting the model to propose synonyms or near-synonyms that could feasibly replace the masked phrase. The MEA stage allows for a nuanced analysis of how the interpretation $Y$ is influenced when $W$ is obscured, yet still impacted by the entities $E$. This stage is instrumental in mitigating the literal interpretation bias of $W$, focusing instead on deriving meaning from the broader linguistic factors present in $X$ and $E$.
@@ 3.2 Context-based Causal Intervention
The ERI stage plays a pivotal role in our FOCUS methodology, where we explore the impact of varying entity relationships on the interpretation of phrases to delve deeper into context-based interpretation. In this stage, we implement an entity replacement strategy combined with a dropout rate, selectively altering entities within the context $X$. This process, denoted as $Y_{ERI},\hat{W}_{ERI}=\mathcal{M}(\tilde{X}_{\text{masked}},\tilde{E}_{\text{replaced}})$, involves generating multiple context sentences with replaced entities, while a dropout rate determines the extent of entity replacement. The addition of this dropout strategy is not only to introduce randomness but also to prevent the emergence of new entity biases that could arise from completely replacing all entities. By maintaining a mix of original and altered entities, we ensure the robustness of understanding while mitigating the risk of introducing new biases. The language model $\mathcal{M}$ then conducts multiple inferences on these varied contexts to synthesize a cohesive interpretation. This approach enhances the robustness of the model by considering a diverse range of entity interactions. The ERI stage is fundamental in understanding the nuanced effects of entity dynamics within different contexts, providing a more comprehensive causal analysis of phrase interpretation.
Model Method F1 Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
@@
Final Synthesis (FS)
The FS stage is the integrative phase of the FOCUS methodology, symbolized as $Y_{FS}=\mathcal{M}(Y_{DI},Y_{MEA},Y_{ERI})$. This stage synthesizes insights from Direct Inquiry (DI), Masked Entity Analysis (MEA), and Entity Replacement Inquiry (ERI). It evaluates the interplay between direct, contextually-driven, and variable entity interpretations to forge a comprehensive understanding of the phrase $W$ within context $X$ and its relation to entities $E$. In FS, we assess causal relationships, reconciling the varied interpretations and confounding factors identified in previous stages. This synthesis refines our understanding of language model interpretations, emphasizing contextual richness and linguistic nuance. The FS stage thus concludes the FOCUS methodology, offering a multi-dimensional perspective on language model analytics.
4 Experiments
4.1 Setup
Large Language Models
Baselines
FOCUS
The outputs from the first three stages were processed by a parser that retained only the final conclusions, filtering out the model’s “reasoning" process. In the fourth stage, the language model was briefed about the concise procedures and objectives of the previous three methods. This briefing included potential key information each method might yield and the risks of information omission. Additionally, to control for syntactic variability and minimize error, the prompts for all baselines, including FOCUS, incorporated the same three sample entries for few-shot in-context learning. This ensured that the language model’s output syntax closely resembled that of the dataset.
Experiment F1 Precision Recall BLUE ROUGE-1 ROUGE-2 ROUGE-L Similarity SimCSE ACC (%)
@@ 4.1 Setup
4.2 Results
Factual dataset
Our study meticulously evaluated the performance of various language models on the factual dataset, with a particular focus on interpreting new phrases. Table 1 summarizes these findings. We observed that the FOCUS consistently outperformed other approaches across all metrics. Specifically, for the GPT-3.5 model, FOCUS achieved the highest F1 score of 0.4292, Precision of 0.4153, and Recall of 0.4524, along with the best BLUE, ROUGE, Similarity, SimCSE scores, and Accuracy at 84.5%. A similar trend was noted with the GPT-4 model. Here, FOCUS again led the metrics, notably achieving an F1 score of 0.4446, Precision of 0.4280, Recall of 0.4714, and an impressive Accuracy of 88.2%. These results underscore FOCUS’s superior ability to interpret and contextualize slang and idiomatic expressions from UrbanDictionary, as compared to other existing methodologies. The comparative analysis clearly illustrates that our FOCUS approach excels in understanding the nuanced and evolving nature of internet language, thereby marking a significant advancement in the field of natural language processing.
Counterfactual dataset
Following the analysis of the factual dataset, we extended our evaluation to the counterfactual dataset, which presents hypothetical scenarios altering real-world language usage. The performance results, as detailed in Table 2, reveal insightful trends. Notably, our FOCUS method again excelled, outperforming others in metrics with GPT-4, achieving an F1 score of 0.4532, Precision of 0.4598, Recall of 0.4551, and an Accuracy of 84.9%. This reinforces FOCUS’s adeptness at interpreting not only direct language use but also hypothetical, contextually modified scenarios. Comparatively, on the factual dataset, while FOCUS led with substantial margins, the gap in performance was more pronounced in the counterfactual dataset. This discrepancy underscores the enhanced complexity in interpreting altered linguistic contexts, a challenge adeptly met by FOCUS. In essence, our results from both datasets affirm the robustness and versatility of FOCUS in navigating the multifaceted nature of language, effectively bridging the gap between real and hypothetical language use.
@@ 4.3 Ablation Study
Our ablation study, focusing on the FOCUS methodology’s components MEA and ERI, utilized GPT-4 on the factual dataset. As shown in Table 3, these components significantly enhance the interpretative capabilities of the model.
MEA (Masked Entity Analysis)
MEA, represented mathematically as $P(Y|W,X,E;\mathcal{M})$, where $Y$ is the interpretation, $W$ the phrase, $X$ the context, and $E$ the entities, critically influences the model’s performance. Its exclusion (w/o MEA) reduced the F1 score to 0.4366 from 0.4446. This result illustrates MEA’s role in disentangling the direct influence of $W$ and $X$ from confounding entities $E$, vital for context-driven interpretation.
ERI (Entity Replacement Inquiry)
The ERI process, which examines the causal links $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$, also shows significant impact. Removing ERI (w/o ERI) decreased the F1 score to 0.4283. ERI’s function in the model, isolating the entity’s influence and exploring alternative causal pathways, proves essential for nuanced language interpretation.
While incorporating either MEA or ERI individually into direct inquiry enhances model performance, their combined use in the FOCUS framework is indispensable for achieving optimal results. This synergy underscores the importance of a comprehensive causal analysis, balancing context and entity dynamics, for the nuanced interpretation of evolving linguistic phenomena.
@@ 5.1 Knowledge Update Methods
LLMs enhance knowledge through Parameter-Efficient Fine-Tuning (PEFT) methods like task-specific parameter addition Houlsby et al. (2019) and low-rank adaptation (LoRA) Hu et al. (2021); Pfeiffer et al. (2020). Despite their efficiency, these methods face challenges such as computational demands, catastrophic forgetting, and reduced task-specific effectiveness Lester et al. (2021). BitFit Ben-Zaken et al. (2021), while simplifying fine-tuning, relies heavily on dataset quality. Trade-offs in computational demands, flexibility, and task compatibility are essential considerations for these methods. Retrieval-Augmented Generation Lewis et al. (2020) and In-context learning-based knowledge editing Zhong et al. (2023) offer dynamic integration of external information, focusing more on fact retrieval than on enhancing deeper understanding.
@@ 6 Conclusion
In this work, we have explored the dynamic and evolving nature of internet language, particularly slang and memes, and their impact on the adaptability of large language models (LLMs). Our study introduced a novel benchmark, SLANG, to assess LLMs’ proficiency in comprehending emerging linguistic trends. Additionally, we proposed the FOCUS methodology, which utilizes causal inference to enhance understanding of new phrases and usage patterns, going beyond traditional models in precision and relevance. Our analysis revealed that the causal inference-based approach of FOCUS significantly outperforms traditional models in interpreting internet slang and memes. This success is underpinned by our rigorous methodologies involving the construction of datasets from UrbanDictionary and the incorporation of both factual and counterfactual instances to provide diverse linguistic contexts. The results from our experiments demonstrate the enhanced capability of LLMs, equipped with our FOCUS method, to adapt to the rapid evolution of online language. This research marks a significant advancement in the field of natural language processing, highlighting the importance of contextual understanding and adaptability in LLMs for effectively navigating the complexities of evolving human communication.
7 Acknowledge

### 20 Feb 2024 v4

@@ 3.2 Context-based Causal Intervention
In the Entity Replacement Inquiry (ERI) stage, we alter entities in context $X$ to assess phrase interpretation variability. We use a dropout rate for entity alteration in $Y_{ERI},\hat{W}_{ERI}=\mathcal{M}(\tilde{X}_{\text{masked}},\tilde{E}_{\text{replaced}})$. This introduces a balance of original and new entities, enhancing model robustness without bias. ERI helps understand entity dynamics’ effects on interpretation, providing deeper causal analysis.
Synthesis (SY)
In the Synthesis (SY) stage, we integrate insights from Direct Inquiry, Masked Entity Inquiry, and Entity Replacement Inquiry. Represented as $Y_{FS}=\mathcal{M}(Y_{DI},Y_{MEA},Y_{ERI})$, this phase evaluates the interplay between direct, contextual, and entity-variable interpretations. SY reconciles varied interpretations and confounders, refining the model’s understanding of language nuances. This final stage offers a multi-dimensional perspective on language model analytics, emphasizing contextual richness.
4 Experiments
4.1 Setup
Large Language Models
Baselines
4.2 Results
Results on factual dataset
Our study meticulously evaluated the performance of various language models on the factual dataset, with a particular focus on interpreting new phrases. Table 1 summarizes these findings. We can see that the FOCUS consistently outperforms other approaches across all metrics. Specifically, for the GPT-3.5 model, FOCUS achieves the highest $\text{F}_{1}$ score of 0.4292, precision of 0.4153, and recall of 0.4524, along with the best BLUE, ROUGE, similarity, SimCSE scores, and accuracy at 84.5%. A similar trend is noted with the GPT-4 model. Here, FOCUS again lead the metrics, notably achieving an $\text{F}_{1}$ score of 0.4446, precision of 0.4280, recall of 0.4714, and an impressive accuracy of 88.2%. These results underscore FOCUS’s superior ability to interpret and contextualize slang and idiomatic expressions from UrbanDictionary, as compared to other existing methodologies.
Model Prompting Method $\textbf{F}_{\textbf{1}}$ Precision Recall BLUE ROUGE Similarity SimCSE ACC (%)
@@ 4.2 Results
Experiment $\textbf{F}_{\textbf{1}}$ Precision Recall BLUE ROUGE-1 ROUGE-2 ROUGE-L Similarity SimCSE ACC (%)
@@
Results on counterfactual dataset
Following the analysis of the factual dataset, we extended our evaluation to the counterfactual dataset, which presents hypothetical scenarios altering real-world language usage. The performance results, as detailed in Table 2, reveal insightful trends. Notably, our FOCUS method again outperforms others in metrics with GPT-4, achieving an $\text{F}_{1}$ score of 0.4532, precision of 0.4598, recall of 0.4551, and an accuracy of 84.9%. This reinforces FOCUS’s adeptness at interpreting not only direct language use but also hypothetical, contextually modified scenarios. In essence, our results from both datasets affirm the robustness and versatility of FOCUS in navigating the multifaceted nature of language, effectively bridging the gap between real and hypothetical language use.
@@ 4.3 Ablation Study
Our ablation study, focusing on the FOCUS methodology’s components MEI and ERI, utilized GPT-4 on the factual dataset. As shown in Table 3, these components significantly enhance the interpretative capabilities of the model.
MEI
The MEI stage, represented mathematically as $P(Y|W,X,E;\mathcal{M})$, where $Y$ is the interpretation, $W$ the phrase, $X$ the context, and $E$ the entities, critically influences the model’s performance. Its exclusion (w/o MEI) reduced the $\text{F}_{1}$ score to 0.4366 from 0.4446. This result illustrates MEA’s role in disentangling the direct influence of $W$ and $X$ from confounding entities $E$, vital for context-driven interpretation.
ERI
The ERI stage, which examines the causal links $X\rightarrow Y\leftarrow E$ and $X\rightarrow E$, also shows significant impact. Removing ERI (w/o ERI) decreased the $\text{F}_{1}$ score to 0.4283. ERI’s function in the model, isolating the entity’s influence and exploring alternative causal pathways, proves essential for nuanced language interpretation.
While incorporating either MEI or ERI individually into direct inquiry enhances model performance, their combined use in the FOCUS framework is indispensable for achieving optimal results. This synergy underscores the importance of a comprehensive causal analysis, balancing context and entity dynamics, for the nuanced interpretation of evolving linguistic phenomena.
@@ 5.1 Knowledge Update Methods
LLMs enhance knowledge through parameter-efficient fine-tuning methods like task-specific parameter addition Houlsby et al. (2019) and low-rank adaptation (LoRA) Hu et al. (2021); Pfeiffer et al. (2020). Despite their efficiency, these methods face challenges such as computational demands, catastrophic forgetting, and reduced task-specific effectiveness Lester et al. (2021). BitFit Ben-Zaken et al. (2021), while simplifying fine-tuning, relies heavily on dataset quality. Trade-offs in computational demands, flexibility, and task compatibility are essential considerations for these methods. retrieval-augmented generation Lewis et al. (2020) and in-context learning-based knowledge editing Zhong et al. (2023) offer dynamic integration of external information, focusing more on fact retrieval than on enhancing deeper understanding.
@@ 6 Conclusion
In this work, we have explored the dynamic and evolving nature of internet language, particularly slang and memes, and their impact on the adaptability of large language models (LLMs). Our study introduced a novel benchmark, SLANG, to assess LLMs’ proficiency in comprehending emerging linguistic trends. Additionally, we proposed the FOCUS methodology, which utilizes causal inference to enhance understanding of new concepts, going beyond other methods in precision and relevance. This success is underpinned by our rigorous methodologies involving the construction of datasets from UrbanDictionary and the incorporation of both factual and counterfactual instances to provide diverse linguistic contexts. The results from our experiments demonstrate the enhanced capability of LLMs, equipped with our FOCUS method, to adapt to the rapid evolution of online language. This research marks a significant advancement in the field of natural language processing, highlighting the importance of contextual understanding and adaptability in LLMs for effectively navigating the complexities of evolving human communication.
Limitations
The main limitation of this work lies in its reliance on Internet-specific data, primarily sourced from UrbanDictionary. While this provides a rich and current linguistic landscape, it may not comprehensively represent the vast diversity of global Internet slang and memes, potentially limiting the generalizability of our findings. Additionally, the performance of FOCUS heavily depends on the quality and variety of the dataset used. Our approach, although innovative, does not fully address the complexities of linguistic evolution in non-English or morphologically rich languages. Future work should explore more diverse datasets and extend our methodology to other languages and linguistic contexts.
Ethics Statement
Our research acknowledges that while methods like the SLANG benchmark and FOCUS approach enhance LLM’s understanding of Internet language, they cannot entirely eliminate the propagation of harmful content. Users must exercise caution and cultural sensitivity, especially when interpreting slang and memes, to avoid reinforcing stereotypes or biases. Our work encourages responsible use, emphasizing the importance of respecting diverse linguistic origins and the natural evolution of language.
Acknowledge